i have a equation/word problem that i can't figure out.
the question is: a 90 inch pipe is cut into two pieces. one piece "x" is 20 inches more than the other piece "y". find the length of each piece.
i came up with 20 inches, and 70 inches. i tried the problem different ways, and can't seem to come up with the correct answer.

Answer by 067isa(13)

the total lenght is 90 . we cut it into two pieces, the firt one is X and the second one is 20 more than the first cut/
that all you need to form an equation of the total length
90= first cut + second cut
90 = X + (X+20)
90=2x+20 " subtruct 20 form both sides "
70=2X divide both by 2
X= 35 which is the first cut
this means the second cut is 20 +35= 55
55+35=90 right